### v1

# A Survey of Generative AI for Intelligent Transportation Systems

Abstract.
Intelligent transportation systems play a crucial role in modern traffic management and optimization, greatly improving traffic efficiency and safety. With the rapid development of generative artificial intelligence (Generative AI) technologies in the fields of image generation and natural language processing, generative AI has also played a crucial role in addressing key issues in intelligent transportation systems, such as data sparsity, difficulty in observing abnormal scenarios, and in modeling data uncertainty. In this review, we systematically investigate the relevant literature on generative AI techniques in addressing key issues in different types of tasks in intelligent transportation systems. First, we introduce the principles of different generative AI techniques, and their potential applications. Then, we classify tasks in intelligent transportation systems into four types: traffic perception, traffic prediction, traffic simulation, and traffic decision-making. We systematically illustrate how generative AI techniques addresses key issues in these four different types of tasks. Finally, we summarize the challenges faced in applying generative AI to intelligent transportation systems, and discuss future research directions based on different application scenarios.
Intelligent transportation system, generative AI, autonomous driving, traffic flow
copyright: acmcopyrightjournalyear: 2022journal: CSURcopyright: none
1. Introduction
The transportation system contains a range of traffic infrastructure, equipment, and management strategies, including roads, traffic control, and traffic planning, designed to manage traffic flow and meet people’s travel needs. Its efficiency significantly impacts both societal and the economic activities. The rapid urbanization has led to a steep increase in the number of vehicles, resulting in prevalent issues like congestion and accidents. These challenges inconvenience travelers and disrupt urban operations, emphasizing the urgency of traffic management and optimization. In recent years, the rapid advancement of computer technology has given rise to Intelligent Transportation Systems (ITS). ITS harnesses cutting-edge technologies, such as artificial intelligence, big data analysis, and the Internet of Things, to create a comprehensive system integrating people, roads, and vehicles. Its goals include reducing congestion, enhancing safety, conserving energy, lowering carbon emissions, and supporting autonomous driving technology.
A fundamental component for realizing ITS is the effective utilization of traffic data. For traffic authorities, they benefit from real-time data collected through sensors, cameras, and various devices, enabling them to monitor and manage road conditions. Furthermore, analyzing historical traffic data can be helpful for predictions regarding future traffic flow and congestion. For travelers, they benefit from real-time traffic data through navigation apps and traffic management systems, which provide up-to-date traffic information to assist in route selection and congestion avoidance. In addition, autonomous vehicles heavily rely on numerous sensors and data to perceive their surroundings, make effective decisions for safe and efficient driving. These aspects indicate the important of traffic data within intelligent transportation systems.
In this context, traditional deep learning methods have been widely adopted due to their efficient application in traffic data analysis and modeling. These algorithms can process massive amounts of traffic data, automatically recognize patterns, predict trends, and optimize decisions, thereby improving traffic efficiency and safety. For example, traditional deep learning techniques such as convolutional neural networks (CNNs) are used to analyze traffic camera images, further identifying vehicles, pedestrians, traffic signs, and road conditions to monitor traffic flow and situations. Recurrent neural networks (RNNs) are suitable for processing sequence data, such as time series data generated by traffic sensors, and can be used to predict traffic flow, congestion, and the likelihood of accidents. Deep reinforcement learning (DRL) methods are applied to the development of autonomous driving technology, training models to make driving decisions by simulating the behavior of intelligent agents (such as autonomous vehicles) in traffic environments. These models can enable safe autonomous driving. By utilizing these deep learning models, intelligent transportation systems can analyze road condition patterns more effectively, predict potential congestion points, and provide adaptive decision support for optimizing traffic efficiency.
However, with the increasing complexity of transportation networks, the interweaving of various modes of transportation, and the growing demand for travel, the traffic environment has become increasingly intricate. This complexity poses numerous challenges for traditional deep learning methods in addressing traffic-related issues.
  • Large amounts of sparse or low-quality data. Traffic data is often sparse and frequently suffers from missing values or noise contamination. Traditional deep learning methods impose high requirements on data quality, especially supervised learning approaches that typically require a substantial amount of labeled input and output data for supervised training. The scarcity or absence of traffic data can significantly impact model performance.
  • Rare abnormal scenarios. Traffic environments exhibit high levels of dynamism, with occasional occurrences of anomalies such as traffic accidents that pose significant threats to traffic safety. These anomalies are challenging to capture as they are often rare events, resulting in a lack of relevant samples in training data. Traditional deep learning methods are more suitable for processing tasks under normal circumstances, and struggle to effectively address rare or previously unseen abnormal situations.
  • Unexplored modeling of uncertainty. In the transportation field, there are many factors that may cause uncertainty. For example, adverse weather conditions such as heavy rain, snow, or smog, as well as the uncertainty of various traffic participants’ behaviors such as changing lanes and emergency braking, make traffic flow prediction more challenging. Traditional deep learning methods usually adopt deterministic approaches, relying on pre-processed features, thus making it difficult to effectively capture such uncertainty.
Generative Artificial Intelligence (Generative AI) technology has recently developed rapidly. It can automatically generate various forms of content based on user input, demands, or instructions, including but not limited to text, images, videos, audio, etc. The emergence of generative AI technology makes content production more efficient and convenient, and has broad application prospects. In the field of generative AI, many mature technologies have been widely applied. For example, Variational Auto-Encoder (VAE) (Kingma and Welling, 2013) is a generative adversarial network that can model and generate complex data. Generative Adversarial Network (GAN) (Goodfellow et al., 2020) improves the quality of generated content through the adversarial process of two neural networks (i.e., generators and discriminators). Normalizing Flow (Rezende and Mohamed, 2015) is a normalization technique used for generating complex data distributions, while Energy-Based Model (LeCun et al., 2006) is a generative model based on energy that can effectively improve the quality of generation. Additionally, Generative Model from Physical Process (Liu et al., 2023c) is a generative model based on physical processes that can be used to generate content with physical constraints. Diffusion Model (Sohl-Dickstein et al., 2015) is a probabilistic method that generates data by adding noise to data gradually to reduce its quality. Generative Pre-trained Transformer (GPT) (Ghojogh and Ghodsi, 2020) is based on the Transformer architecture (Vaswani et al., 2017), which pre-trains a large amount of text data and then uses it to generate new text content.
These generative AI techniques offer significant advantages in data generation, transformation, and repair, and they can effectively model the uncertainty of data. As a result, these technologies have garnered significant attention in different fields. Especially in the field of intelligent transportation, generative AI technology has played a crucial role and introduced innovative solutions to address traffic-related challenges (Lv, 2023). Numerous researchers are actively engaged in the application of diverse generative AI techniques to traffic-related tasks. For example, in autonomous driving scenarios, generative AI technology proves invaluable in generating high-fidelity driving scene images and videos, which are essential for the training and testing of autonomous driving systems (Krajewski et al., 2018; Kumar et al., 2023). Utilizing generative AI-generated simulation scenarios, autonomous vehicles can conduct extensive driving simulations within virtual environments, thereby enhancing their ability to make precise real-world decisions. In traffic flow prediction tasks, generative AI technology can learn and model the distribution of traffic data, thus generating future traffic flow data (Zhang et al., 2022b; Zand et al., 2023; Wen et al., 2023). This is very helpful for traffic authorities to carry out effective traffic planning and optimization.
Category Reference Description
ITS Comprehensive
(Qureshi and Abdullah, 2013)(An et al., 2011)(Singh and Gupta, 2015)(Ghosh et al., 2017)(Figueiredo et al., 2001)(Fayaz, 2018)(Zear et al., 2016)(Garg and Kaur, 2023)(Xiong et al., 2012)(Mandal et al., 2022)(Tran et al., 2023)(Rhoades and Conrad, 2017),(Sharma and Awasthi, 2022)(Hassan et al., 2023)(Ravi and Mamdikar, 2022)
These works give a comprehensive review of ITS, covering
key technologies, infrastructure, and applications. However,
few generative AI techniques are systematically introduced.
Technologies
DRL (Haydari and Yılmaz, 2020), GNN (Rahmani et al., 2023), NLP (Putri et al., 2021), SVM (Hao et al., 2020), DL (Veres and Moussa, 2019; Haghighat et al., 2020), Positioning (Du et al., 2021; Imparato et al., 2018), GAN (Lin et al., 2023a),
Blockchain (Swain, 2006; Das et al., 2023), Edge computing (Zhou et al., 2021a; Gong et al., 2023), Dynamic pricing (Saharan et al., 2020), Fuzzy logic (Swain, 2006),
Swarm intelligence (Mathew, 2020), V2X downloading (Wang et al., 2019), Data mining (Anand et al., 2018; El Faouzi et al., 2011; Zhu et al., 2018),
Communication (Deng et al., 2023; Camacho et al., 2018; Maimaris and Papageorgiou, 2016; Naja, 2013; Shaaban et al., 2021; Marè et al., 2016; Khan et al., 2022)
These works conduct surveys on specific techniques within
the field of in ITS.
Infrastructure
IoT (Chand and Karthikeyan, 2018; Patel et al., 2019; Bhardwaj et al., 2019), VANET (Baras et al., 2018; Trivedi et al., 2019), Bluetooth (Friesen and McLeod, 2015), Radar-on-chip (Saponara et al., 2019), Hardware devices (Damaj et al., 2022),
Distributed architecture (Nasim and Kassler, 2012), External infrastructure (Creß et al., 2021)
These works conduct surveys on specific components in ITS.
Applications
Security management (Ali et al., 2018; Ben Hamida et al., 2015; Lamssaggad et al., 2021; Harvey and Kumar, 2020; Amin et al., 2019; Affia et al., 2019), Traffic scheduling (Nama et al., 2021), Traffic detection (Lin et al., 2022),
Driver behavior detection (Chhabra et al., 2017), Vehicle classification (Gholamhosseinian and Seitz, 2021; Yang and Pun-Cheng, 2018), Urban mobility (Mangiaracina et al., 2017),
Urban monitoring (Lima et al., 2017), Traffic sign detection (Ellahyani et al., 2021; Luo et al., 2020), Traffic prediction (Suhas et al., 2017; Yuan and Li, 2021),
Autonomous driving (Altaf and Kaul, 2021), Metaverse (Njoku et al., 2023), Trust management (Ma et al., 2011), Spectrum regulation (Choi et al., 2020)
These surveys focus on specified applications in ITS.
Generative AI Comprehensive
(Cao et al., 2023)(Wu et al., 2023)(Zhang et al., 2023f)(Foo et al., 2023)
These works conduct a comprehensive review of generative
AI techniques.
Applications
3D (Li et al., 2023d), Biology (Zhang et al., 2023b), ChatGPT (Zhang et al., 2023e; Wang et al., 2023a), Metaverse (Qin and Hui, 2023), Edge cloud computing (Wang et al., 2023b),
Mobile network (Xu et al., 2023a; Du et al., 2023), Security & Privacy (Wang et al., 2023c; Chen et al., 2023b), Vehicle network (Zhang et al., 2023d)
These surveys concentrate on the specific applications of
generative AI techniques in a particular field.
Table 1. Related reviews of intelligent transportation system (ITS) and generative AI. DRL: Deep reinforcement learning, GNN: Graph neural network, NLP: Natural language processing, V2X: Vehicle to everything. DL: Deep learning. IoT: Internet of things. VANET: Vehicle ad-hoc network.
As shown in Table 1, there are numerous surveys on intelligent transportation systems or generative AI. However, most of them rarely explore the practical applications of generative AI technology in intelligent transportation systems. Although the authors in (Lin et al., 2023a) have explored the role of GANs in intelligent transportation systems, they do not comprehensively cover other generative AI techniques. Importantly, they do not systematically examine the advantages of generative AI techniques over traditional deep learning methods in intelligent transportation systems. Therefore, in this paper, we will explore the applications of generative AI techniques in traffic perception, prediction, simulation, and decision-making within intelligent transportation systems, as shown in Figure 1. Our goal is to provide a comprehensive analysis of their role in promoting intelligent transportation, and offer a constructive insight to help readers gain a deeper understanding of the challenges and opportunities of generative AI in intelligent transportation systems. We expect this survey to provide useful insights for further research and advancement of intelligent transportation systems.
Figure 1. An overview of generative AI in intelligent transportation systems.
In conclusion, the primary contributions of this paper are illustrated as follows:
  • To the best of our knowledge, we are the first to offer an in-depth literature review of generative AI for intelligent transportation systems.
  • We provide a systematic introduction to mainstream generative AI techniques, conduct in-depth method comparisons from both horizontal and vertical perspectives, and provide a systematical analysis of how generative AI technology can effectively address key issues in intelligent transportation systems from the aspects of traffic perception, traffic prediction, traffic simulation, and traffic decision-making.
  • We discuss the open challenges encountered in applying generative AI technology in intelligent transportation systems, and explore potential directions for future research.
The structure of this paper is as follows. Section 2 introduces mainstream generative AI techniques and their application achievements in different fields, followed by a detailed comparative analysis. The following four sections respectively discuss the applications of generative AI technology in traffic perception, traffic prediction, traffic simulation, and traffic decision-making. Finally, in Section 7, we introduce the challenges faced by generative AI in the application of intelligent transportation systems, and look forward to future research directions.
2. Technologies, Applications and Advantages of Generative AI
As described in previous works (Cao et al., 2023; Foo et al., 2023; Wang et al., 2023a; Wu et al., 2023), generative AI refers to a subset of AI algorithms that focus on creating or producing personalized, and high-quality content, including text, images, videos, and 3D assets, in response to user input or specific requirements. This section provides a comprehensive overview of the underlying technologies, highlights their significant applications, and summarizes the advantages of generative AI techniques in comparison to traditional deep learning methods.
2.1. Key Technologies
With the rapid growth of computer technology and computing power, generative AI technology has witnessed significant progress over the past decade. Variational Autoencoders (VAE), a probabilistic graphical model, was pioneered by Kingma and Welling in 2013 for data generation (Kingma and Welling, 2013). In 2014, Ian Goodfellow et al. designed generative adversarial networks (GANs) (Goodfellow et al., 2014), which led to a major transformation in the field of generative AI. In 2018, OpenAI proposed Generative Pre-trained Transformer (GPT), which also brought important breakthroughs to the field. Beyond these approaches, technologies like Normalizing Flow, Energy-Based Models (EBMs), Diffusion Probabilistic Models (DPMs) and Generative Models from Physical Process (GenPhys) have also demonstrated substantial potential in the domains of text and image generation. Figure 2 depicts the evolution timeline of generative AI techniques, including VAE, GANs, Normalizing Flow, EBMs, DPMs, GenPhys, and GPT. In this subsection, we will provide an initial introduction to these techniques, followed by a comparative analysis.
Figure 2. The history of generative AI techniques.
2.1.1. VAE
VAE (Kingma and Welling, 2013) aims to learn the latent representations of data and generate new data samples. It achieves this by mapping data to a probability distribution in a latent space and then re-mapping the latent variables back to the data space through a decoder.
[…]
2.1.2. GAN
GAN (Goodfellow et al., 2014) consists of a generator and a discriminator. They are trained together through an adversarial process to generate realistic data samples. The generator $G$ takes random noise as input, drawn from a prior distribution $p_{z}(z)$, and generates fake data samples $G(z)$ that resemble real data samples. The discriminator $D$ receives both real data samples $x$ from the dataset following $p_{\text{data }}(\boldsymbol{x})$ and fake data samples $G(z)$ generated by the generator. Its objective is to differentiate between real and fake samples. The objective of GAN is to find a Nash equilibrium where the generator generates data that is indistinguishable from real data. This equilibrium is achieved when the discriminator cannot reliably distinguish between real and fake samples.
2.1.3. Normalizing Flows
The core idea of normalizing flows (Rezende and Mohamed, 2015) is to map a simple probability distribution to a complex one by applying a sequence of reversible transformations, so as to generate samples that match the complex data distribution. These reversible transformations can be combined using the chain rule to efficiently compute the logarithmic probability density of the generated samples. The key feature is that these transformations are all reversible, allowing for easy mapping from the data space to the latent space and vice versa.
Let $f$ be an invertible and smooth mapping function, and its inverse function $f^{-1}=g$. By transforming a random variable $\mathbf{z}\sim p(\mathbf{z})$ using $f$, the distribution of the resulting random variable $\mathbf{y}=f(\mathbf{z})$ is:
(1)1( 1 ) $p\left(\mathbf{y}\right)=p(\mathbf{z})\left|\operatorname{det}\frac{\partial f^{-1}}{\partial\mathbf{y}}\right|=p(\mathbf{z})\left|\operatorname{det}\frac{\partial f}{\partial\mathbf{z}}\right|^{-1},$
where the chain rule is applied. By combining multiple simple mappings and successively using Equation 1, arbitrarily complex densities can be derived. The density $p_{K}(\mathbf{z})$ calculated by the successive transformations of a random variable $\mathbf{z}_{0}$ following the distribution $p_{0}$ through a chain of $K$ transformations is written as:
(2)2( 2 ) $\mathbf{z}_{K}=f_{K}\circ\ldots\circ f_{2}\circ f_{1}\left(\mathbf{z}_{0}\right),$
(3)3( 3 ) $\ln p_{K}\left(\mathbf{z}_{K}\right)=\ln p_{0}\left(\mathbf{z}_{0}\right)-\sum_{k=1}^{K}\ln\left|\operatorname{det}\frac{\partial f_{k}}{\partial\mathbf{z}_{k-1}}\right|.$
Here, the path that is traversed by the random variables $\mathbf{z}_{k}=f_{k}(\mathbf{z}_{k-1})$, which are generated from the initial distribution $p_{0}(\mathbf{z}_{0})$, is known as the flow. The path created by the consecutive distributions $p_{k}$ is referred to as a normalizing flow. By selecting suitable transformations $f_{k}$, we can begin with simple, factored distributions like an independent Gaussian. Then, by applying normalizing flows of varying lengths, we can obtain increasingly complex and multi-modal distributions.
2.1.4. Energy-Based Models
EBMs (LeCun et al., 2006; Du and Mordatch, 2019) focus on defining an energy function that assigns a scalar value to each data point. The energy function measures how well a data point fits the model, with lower energy values indicating a better fit. EBMs aim to assign lower energy values to realistic data points and higher energy values to data points that are not part of the true distribution. In other words, they can approximate the likelihood of the observed data and offer a quantitative assessment of how well a given data point aligns with the learned distribution.
For the observed data $\mathbf{x}$, its probability density function $p_{\theta}(\mathbf{x})$ is defined via the Boltzmann distribution:
(4)4( 4 ) $p_{\theta}(\mathbf{x})=\frac{e^{-E_{\theta}(\boldsymbol{\mathbf{x}})}}{\int_{\tilde{\mathbf{x}}\in\mathcal{X}}e^{-E_{\theta}(\tilde{\mathbf{x}})}},$
where $E_{\theta}(\mathbf{x})$ is an energy function with parameters $\theta$.
A common approach to optimize EBMs is contrastive divergence. In contrastive divergence, the model first samples a sample from the positive data distribution and another sample from the negative data distribution. Subsequently, the model updates its parameters by comparing the energy values of these two samples. The objective is to decrease the energy value of positive samples while increasing the energy value of negative samples. This process is typically repeated iteratively to gradually optimize the model.
2.1.5. Diffusion Probabilistic Model
The essence of DPMs (Sohl-Dickstein et al., 2015) lies in its step-by-step denoising process that gradually recovers the underlying original data samples.
[…]
In the forward process, a Markov chain successively adds Gaussian noise to the data $\mathbf{x}_{0}$. Hence, we can obtain the approximate posterior $q\left(\boldsymbol{x}_{1:T}\mid\boldsymbol{x}_{0}\right)$:
(5)5( 5 ) $q\left(\boldsymbol{x}_{1:T}\mid\boldsymbol{x}_{0}\right)=\prod_{t=1}^{T}q\left(\boldsymbol{x}_{t}\mid\boldsymbol{x}_{t-1}\right),$
where $q\left(\boldsymbol{x}_{t}\mid\boldsymbol{x}_{t-1}\right)$ is a Gaussian distribution.
In the reverse process, the original data point is recovered by iteratively removing noise from the noisy data. Specifically, this process is structured as a Markov chain, leveraging learnable Gaussian transitions that initiate from the distribution $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};\mathbf{0},\mathbf{I})$:
(6)6( 6 ) $p_{\theta}\left(\boldsymbol{x}_{0:T}\right)=p\left(\boldsymbol{x}_{T}\right)\prod_{t=T}^{1}p_{\theta}\left(\boldsymbol{x}_{t-1}\mid\boldsymbol{x}_{t}\right),$
where $p_{\theta}(\boldsymbol{x}_{t-1}\mid\boldsymbol{x}_{t})$ represents a reverse process.
During each training iteration, the network attempts to predict the exact noise added to the data at a specific diffusion step. The training objective is to minimize the difference between the network’s predicted noise and the actual noise.
2.1.6. Generative Models from Physical Process
GenPhys (Liu et al., 2023c) transform physical partial differential equations (PDEs) into generative models. They provide a more general generative framework, with diffusion probabilistic model being a special case within the GenPhys framework.
Given observed data sampled from the probability distribution $p_{data}(\mathbf{x})$, the task aims to produce new sample from $p_{data}(\mathbf{x})$. A continuous physical process can be described by
(7)7( 7 ) $\frac{d\mathbf{x}}{dt}=\mathbf{v}(\mathbf{x},t),$
where $\mathbf{v}$ represents a velocity field.
This process dynamically evolves the probability distribution $p(\mathbf{x},t)$ according to the following equation:
(8)8( 8 ) $\frac{\partial p(\mathbf{x},t)}{\partial t}+\nabla\cdot[p(\mathbf{x},t)\mathbf{v}(\mathbf{x},t)]-R(\mathbf{x},t)=0,$
where $R(\mathbf{x},t)$ can be interpreted as birth or death. When $R>0$, it means indicates the birth of particles in the forward process and their subsequent death in the backward process, and vice versa. This equation is known as density flow equation. The goal is to design $p(\mathbf{x},t),\mathbf{v}(\mathbf{x},t)$ and $R(\mathbf{x},t)$ in a manner that fulfills two conditions. First, it should ensure that $p(\mathbf{x},0)=p_{data}(\mathbf{x})$. The second condition is that $p_{\text{prior}}(\mathbf{x})\equiv p(\mathbf{x},T)$ becomes asymptotically independent of $p_{\text{data}}(\mathbf{x})$ as $T\rightarrow\infty$.
Continuous physical processes are governed by PDEs, which is written as:
(9)9( 9 ) $\hat{L}\phi\equiv F\left(\phi,\phi_{t},\phi_{tt},\nabla\phi,\nabla^{2}\phi,\ldots\right)=f(\mathbf{x},t),$
where $\hat{L}$ represents a differential operator on the scalar function $\phi(\mathbf{x},t)$. $f(\mathbf{x},t)$ is the source term.
Next, the crucial step is to map the solutions of physical PDEs to those of density flows. This requires density flows and physical PDEs are equivalent equations. The final step is to produce $\mathbf{x}(0)$ by the reverse physical process. To begin, $\mathbf{x}(T)$ is drawn from the distribution $p(\mathbf{x},T)$. Then, we simulate the backward evolution of $\frac{d\mathbf{x}}{dt}=\mathbf{s}_{\theta}(\mathbf{x},t)$ from $t=T$ to $t=0$ using the branching process $W_{\alpha}(\mathbf{x},t)$. This simulation ultimately produces $\mathbf{x}(0)$.
2.1.7. Generative Pre-trained Transformer
Generative Pre-trained Transformer (GPT) is a neural network framework based on Transformer, specifically developed for tasks involving natural language processing (NLP) and generation. GPT models are pre-trained on a large number of textual datasets and are capable of generating text that is both coherent and contextually meaningful. GPT utilizes a transformer architecture (Vaswani et al., 2017) with decoder layers. Each decoder layer consists of self-attention mechanisms and feedforward neural networks. These mechanisms allow the model to capture long-range dependencies in sequential data like text. To handle sequential input data, GPT includes positional encodings and learned word embeddings. Layer normalization is applied after each sub-layer in the transformer decoder to stabilize training. GPT employs an autoregressive generation approach, producing output tokens sequentially, building upon previously produced tokens. An important feature of GPT is parameter sharing across its decoder layers, ensuring consistency in data representation. GPT models undergo unsupervised pre-training on extensive text corpora and can be fine-tuned for specific tasks with labeled data. Their scalability has led to increasingly large models, such as GPT-2 (Nguyen et al., 2021), GPT-3 (Brown et al., 2020) and GPT-4 (OpenAI, 2023), achieving superior performance across various NLP tasks.
2.1.8. Comparative Analysis of Generative AI Techniques
In previous subsections, we have introduced several major generative AI techniques. Each of them has its distinct advantages and limitations. For example, VAE offer a probabilistic framework for encoding data into a lower-dimensional space but often face challenges in selecting the appropriate likelihood functions and generating high-quality images. GANs aim to generate highly realistic data samples by pitting a generator against a discriminator in a adversarial training scheme. They produce impressive results in image generation but can be challenging to train and prone to mode collapse. We summarize the advantages and disadvantages of these generative AI techniques, as shown in Table 2.
Technologies Advantages Disadvantages
Variational Autoencoder (Kingma and Welling, 2013)
- Explicit encoder and decoder
- Capable of probabilistic generation, suitable for data with uncertainty
- Can learn latent representations of data
- Training is relatively stable
- May produce blurry images
- Challenges in selecting likelihood functions
- Limited expressiveness of the prior distribution
Generative Adversarial Network (Goodfellow et al., 2020)
- Capable of generating high-quality images
- Provides diversity in the generation process
- Training process can be unstable
- Mode collapse issue is common
Normalizing Flow (Rezende and Mohamed, 2015)
- Can accurately model probability distributions
- Supports both generation and inference
- Applicable to various data types
- Training and inference complexities are high
- Requires more parameters for high-dimensional data
Energy-Based Model (LeCun et al., 2006)
- Can model complex probability distributions
- Supports both generation and inference
- Robust to outliers
- Requires careful tuning of hyperparameters
- Training process may be slow
Diffusion Probabilistic Model (Sohl-Dickstein et al., 2015)
- Robust to noise and uncertainty
- Flexible noise schedule choice
- Allows for controlled generation
- Longer generation time
- Training complex diffusion models can be slow
Generative Models from Physical Process (Liu et al., 2023c)
- Generates data based on physical processes
- Offers interpretability
- Requires prior knowledge of physical parameters
- Typically limited to specific application domains
Generative Pre-trained Transformer (Ghojogh and Ghodsi, 2020)
- Performs exceptionally well in natural language processing tasks
- Capable of generating continuous text
- Effective modeling of long-term dependencies
- Requires substantial computing resources for training
- Dependence on large training datasets
Table 2. The advantages and disadvantages of Generative AI techniques, partly referenced by the work (Zhang et al., 2023d).
2.2. Important Generation Tasks
The advancement of generative technologies has provided impressive solutions for various generative tasks. In this section, we will explore the applications of generative technologies in these generation tasks from the perspectives of text, images, videos, and cross-modal data.
2.2.1. Text Generation
Text generation focuses on using AI techniques to autonomously generate reasonable and meaningful text. This technology is often used in tasks such as machine translation, chatbots, and article summarization. In recent years, with the advancement of deep learning technologies, especially the emergence of the Transformer architecture and the GPT series of models, both the quality and diversity of text generation have significantly improved. ChatGPT (Ouyang et al., 2022), developed by OpenAI, is a specialized version of the GPT model fine-tuned for conversational interactions. Microsoft’s Bing (Adiwardana et al., 2020) has integrated ChatGPT technology, enhancing the conversational search experience for users. This integration empowers users to comprehend intricate queries, receive detailed responses, and benefits from continuous learning and optimization capabilities.
2.2.2. Image Generation
Image generation focuses on using AI algorithms to create new images or visual content using given information such as images and text. For example, generative methods leverage the power of neural networks to learn complex data representations, making them a prominent approach for restoring images to their original high-quality state (Sohl-Dickstein et al., 2015; Ledig et al., 2017; Cai and Wei, 2020; Rombach et al., 2022). Several works focus on conversion between two images, encompassing tasks such as style transfer and image morphing (Karras et al., 2019; Zhao et al., 2021b). Recently, diffusion models are utilized to address the tasks of image editing, achieving superior image generation quality (Li et al., 2022e; Ackermann and Li, 2022; Avrahami et al., 2023).
2.2.3. Video Generation
Video generation involves creating a series of frames that are semantically coherent and appear in a sequence to form a video. This task is challenging due to its requirement to deal with high-dimensional video data and simultaneously model spatial and temporal depenedencies. The related works are usually classified into two types. Unconditional video generation aims to generate new videos without requiring additional information. It achieves this by extracting valuable information from the training data. Numerous works that tackle the task of unconditional video generation utilize generative AI technology (Wang et al., 2020; Gupta et al., 2021; Wang et al., 2021a). Unlike unconditional video generation, conditional video generation incorporates additional condition to guide the process of video generation. VAE and GAN are are two popular methods used to generate videos with specific settings such as text (Li et al., 2018a), sentences (Liu et al., 2019) and captions (Pan et al., 2017).
2.2.4. Cross-modality Data Generation
Cross-modality data generation refers to generate data in one modality such as image, text and video based on the features from a different modality. The primary challenge in implementing cross-modality data generation lies in effectively modeling the relationships between different modalities, enabling control over the generation process through inputs from another modality (Foo et al., 2023). Text-to-image generation aims to generate images based on a provided textual description. Current approaches for text-to-image generation are mainly built upon GANs (Xu et al., 2018; Li et al., 2019), Transformer (Ramesh et al., 2021; Ding et al., 2021, 2022) and diffusion models (Nichol et al., 2021; Radford et al., 2021). Image-to-text generation involves generating textual description or captions that describe the content of an image. Most works adopt the encoder-decoder architectures to deal with this task. The language decoder plays a crucial role in model design. Advancements in the NLP field have led to a transition in language decoder architecture from RNNs to Transformer-based generative models (Li et al., 2020c; Zhang et al., 2021b). Text-to-video generation refers to creating videos from the textual prompts. Three commonly used approaches to address the challenge of text-to-video generation include GAN-based (Balaji et al., 2019) and VAE-based (Mittal et al., 2017), and diffusion-based (Singer et al., 2022; Rahman et al., 2023) methods.
2.3. Advantages of generative AI technology
Generative AI technology has significant advantages over traditional deep learning methods such as MLP, CNN, RNN in many aspects. In this subsection, we will discuss them in terms of data generation, creativity, data utilization, model generalization, and cross-domain applications.
Data generation. Generative AI not only has the capability to process and analyze existing data but also to generate new data samples based on the distribution of the available data. For instance, in the field of computer vision, GANs can produce highly realistic images. In natural language processing, generative pre-trained models like the GPT series are proficient at generating fluent text, which can be used for tasks such as automated content generation, chatbot systems, and translation services. Furthermore, generative AI finds applications in various domains, including audio synthesis, video generation, medical image synthesis, and simulating experiments in scientific research. While traditional deep learning methods excel in data analysis and classification tasks, they typically lack the direct capability to generate entirely new data samples. This makes generative AI particularly advantageous in data augmentation and simulation among other areas.
Creativity. Generative AI has the ability to generate novel, interesting, and diverse data, which enables it to have extensive application potential in the creative field. For example, in the field of literature, generative pre-trained models such as GPT-3 (Floridi and Chiriatti, 2020) have been able to automatically generate articles, essays, poems, and stories. Generative AI is also used in music composition (Muhamed et al., 2021; Zhang, 2020; Shih et al., 2022). Music generation models can generate various styles and types of music, from classical to popular. In the field of scientific research, generative AI is also used for new drug discovery (Cheng et al., 2021b; Bilodeau et al., 2022). By simulating molecular structures and chemical reactions, generative AI can generate potential new drug candidates, accelerating the process of drug development. Traditional deep learning methods rely more on existing data for extraction and classification, and are relatively weak in terms of creativity.
Reasoning. Generative AI methods have more advantages than traditional deep learning methods in terms of reasoning. Firstly, they often employ probabilistic modeling, which can effectively model uncertainties. Secondly, they have the ability to capture long-term dependencies. For example, the self-attention mechanism in the GPT architecture allows the model to capture information from any position in the sequence during the processing of sequence data, rather than just local dependencies. Traditional deep learning methods typically focus on end-to-end learning, with the main goal of mapping input data to output data rather than specific reasoning. Therefore, the probabilistic modeling and long-term dependency extraction of the generative AI methods make it more outstanding in reasoning.
Model generalization. Generative AI not only learns broader knowledge through large-scale data training, but also benefits from its architecture with numerous shared parameters, allowing it to transfer learned knowledge across different tasks. For example, GPT can be applied to a variety of NLP tasks including text generation, language translation, and dialogue generation, without the need to design new models for each task separately. Similarly, in computer vision, generative AI-based models can initially be trained for image classification, and then are employed for other tasks such as object detection without the requirement for redesigning network structures. Achieving such adaptability and versatility proves challenging for traditional deep learning methods, which typically require task-specific network architectures.
Cross-domain applications. Generative AI methods can enable cross-domain data generation, allowing knowledge acquired in one domain to seamlessly transfer to another. For example, Google’s Imagen uses a diffusion model to convert text descriptions into images (Ho et al., 2022). This cross-domain transfer learning provides new opportunity for knowledge sharing between different domains. In contrast, traditional deep learning methods are relatively weak in cross-domain applications. They usually require redesigning and training models specific to the new domain to adapt to the data and tasks of different domains. This leads to increased complexity and resource requirements for cross-domain knowledge transfer, limiting their extensive applicability across different domains.
3. Generative AI for Traffic Perception
Traffic perception refers to the ability of intelligent transportation systems to collect and understand various sensory information in the traffic environment. This information includes visual data from cameras, trajectory data from GPS-based devices, motion data from accelerometers, environmental data like weather and road conditions, and other traffic-related information such as the presence of other vehicles, pedestrians, and road signs. It plays a crucial role in making effective decisions and ensuring safe driving. In the context of autonomous vehicles, traffic perception typically involves the use of cameras, LiDAR, millimeter-wave radar, and other technologies to collect and process data from the vehicle’s surroundings, enabling autonomous driving and navigation.
However, the accurate perception and comprehension of the intricate dynamics within traffic environments pose several significant challenges: First, sensors may fail to provide data in certain situations, such as when cameras are obstructed. These data gaps can disrupt the continuous monitoring of traffic conditions, potentially leading to incomplete information. Second, sensor data collected in the traffic environment is subject to various sources of noise, including sensor inaccuracies, weather conditions, and lighting variations. This noise introduces uncertainty into the data, making it harder to derive precise information. Third, traffic environments are inherently complex and dynamic, involving numerous interacting factors such as vehicles, pedestrians, traffic signals, and road conditions. This complexity makes it challenging to accurately model and understand traffic behavior.
The advance of generative AI technology holds the potential to offer solutions to these challenges. We carried an in-depth literature review focusing on the application of generative AI to address challenges in traffic perception. The reviewed literature covers various aspects of traffic perception, including traffic data imputation, traffic estimation, traffic data analysis and traffic anomaly detection. Their relation is illustrated in Figure 3.
Figure 3. The relation between different topics in generative AI for traffic perception research.
3.1. Traffic Data Imputation
The task of traffic data imputation involves learning a function, denoted as $f(\cdot)$, which takes as input the corrupted data or missing data $X$, in order to reconstruct the complete data $\hat{X}$. Formally, the reconstructed data $\hat{X}$ can be expressed as $\hat{X}=f(X,\Psi)$, where $\Psi$ denotes other important factors like road network graph.
Low-dimensional traffic data. The low-dimensional traffic data like speed, flow and travel time is crucial for various traffic applications such as route planning and traffic prediction. Missing it is a common issue due to sensor device damage or transmission disruptions. This has led researchers to employ advanced generative AI techniques for imputing missing data using available measurements.
VAE. In previous research (Boquet et al., 2019), a VAE was employed to facilitate the imputation of missing traffic data in an online unsupervised manner by drawing from the learned data distribution. Chen et al. (Chen et al., 2021d) proposed a spatio-temporal VAE method for imputing missing data in traffic raster datasets. This method leverages 3D gated convolution and a multi-attention network to effectively capture spatio-temporal dependencies under the VAE architecture.
GAN. Recently, more researchers have adopted GAN-based methods to address this task (Chen et al., 2019; Han et al., 2020b; Kazemi and Meidani, 2021). For example, Chen et al. (Chen et al., 2019) proposed using real or corrupted data as latent codes and introducing representation loss within a GAN model, significantly improving the performance of traffic data imputation. IGANI (Kazemi and Meidani, 2021) introduced a novel iterative GAN imputation architecture, which iterated over imputed data while preserving the reversibility of the generative imputer.
An important characteristic of traffic data lies in its spatiotemporal correlation. Thus, numerous works proposed to incorporate the spatio-temporal modeling in GAN-based architectures (Huang et al., 2021; Yang et al., 2021; Zhang et al., 2021f, d, d; Yang et al., 2022; Huang and Chen, 2022; Shen et al., 2022; Li et al., 2023c, a; Peng et al., 2023). For example, The authors in (Li et al., 2018b) introduced a 3D convolutional GAN for imputing missing traffic data, utilizing a fractional-strided 3D convolutional neural network (CNN) for the generator and a 3D CNN for the discriminator to capture spatio-temporal features efficiently. STGAN (Yuan et al., 2022) incorporated generation and center losses to minimize reconstruction errors and maintain local spatio-temporal distributions, while the discriminator used a CNN classifier to assess global spatio-temporal distribution compliance, and the generator employed skip connections and dilated convolutions to capture spatio-temporal correlations in traffic data. GAE-GAN-LSTM (Xu et al., 2022) leveraged an enhanced graph autoencoder to extract spatio-temporal features, and employed a GAN for generating complete spatio-temporal features from the missing data, with a generator built on long short-term memory and a discriminator using a fully connected neural network.
Diffusion model. Recently, diffusion models have attracted more researchers’ attentions in traffic data imputation (Liu et al., 2023b; Zhang et al., 2023c; Yun et al., 2023). For instance, Liu et al. (Liu et al., 2023b) introduced an enhanced prior modeling conditional diffusion framework for spatio-temporal imputation, which effectively handles various missing patterns and scenarios, including high missing rates and sensor failures, by leveraging conditional feature extraction and noise estimation modules. SaSDim (Zhang et al., 2023c) introduced a self-adaptive noise scaling mechanism based on diffusion model. This model incorporated a new loss function to scale the noise to a consistent intensity level, and a spatial-temporal global convolution module to capture complex spatial-temporal patterns.
Method Spatio-Temporal Modeling Papers
VAE MLP (Boquet et al., 2019)
3D gated CNN, multi-attention mechanism (Chen et al., 2021d)
GAN 3D CNN (Li et al., 2018b)
Dilated CNN (Yuan et al., 2022)
CNN (Yang et al., 2022; Zhang et al., 2021a)
CNN, self-attention mechanism (Zhang et al., 2021f)
Temporal: Gramian Angular Summation Field (Huang et al., 2021)
Temporal: discrete Wavelet transform (Huang and Chen, 2022)
Temporal: multi-view temporal factorizations (Li et al., 2023c)
Spatial: learnable bidirectional attention map; Temporal: multi-channel matrix
(Yang et al., 2021)
Spatial: GCN; Temporal: LSTM
(Xu et al., 2022; Shen et al., 2022)
Spatial: multi-layer FCN; Temporal: LSTM
(Li et al., 2023a)
Spatial: dynamic GCN; Temporal: multi-head self-attention network
(Peng et al., 2023)
Spatial: spatial attention GCN; Temporal: self-attention GRU
(Zhang et al., 2021d)
DPM Spatial: GNN (Yun et al., 2023)
Spatial: GCN; Temporal: Transformer
(Liu et al., 2023b)
Spatial: dynamic GCN; Temporal: global temporal convolution
(Zhang et al., 2023c)
Table 3. Spatio-temporal modeling in generative AI methods for low-dimensional traffic data imputation.
Finally, to provide a clear illustration of the spatio-temporal modeling within the generative AI framework for this task, we summarize the related works in Table 3.
High-dimensional traffic data. High-dimensional traffic data, such as traffic videos and 3D LiDAR vision, plays a critical role, particularly in the context of autonomous driving, where complete dataset would greatly enhance situational awareness and decision-making capabilities. However, the loss of high-dimensional traffic data remains a prevalent issue in real-world scenarios. To combat this challenge, Wu et al. (Wu et al., 2022) introduced a novel complementary model based on GANs for addressing missing frames in traffic videos. The model utilized a feature pyramid network to extract multi-scale feature maps from input video frames, enhancing the integration of semantic information across frames by fusing features from different scales. The inclusion of local block discriminators within the discriminator model effectively ensured the accuracy and continuity of the generated frames. TDC-GAN (Wu et al., 2021) utilized a multi-scale semantic information extraction model based on the feature pyramid network and a discriminator model with global and local branches to ensure spatio-temporal consistency in consecutive frames. For incomplete point clouds in LiDAR applications, Tu et al. (Tu et al., 2021) introduced a GAN-based point cloud repair network specifically applied within the autonomous driving context, enhancing overall safety in autonomous driving systems.
3.2. Traffic Estimation
Traffic estimation aims to learn a estimator to estimate the current traffic condition of all roads in the transportation network. This task differs from traffic data imputation in that it addresses the challenge of incomplete data collection due to constraints such as cost and time limitations.
Several works designed GAN-based approaches to estimate traffic condition (Liang et al., 2018; Yu and Gu, 2019; Zhang et al., 2021c; Mo et al., 2022; Tian et al., 2022; Zhang et al., 2022c; Xu et al., 2023c). GAA (Liang et al., 2018) incorporated traffic-flow theory into a deep GAN model, while GCGA (Yu and Gu, 2019) combined the graph modeling capabilities of GCN and the generative capabilities of GAN . Curb-GAN (Zhang et al., 2020b) enhanced the conditional GAN structure by modeling various travel demands. GE-GAN (Xu et al., 2020b) learns road network features through graph embedding and applies them to a GAN network to estimate traffic conditions. PA-GAN (Tian et al., 2022) incorporated Bayesian inference to adaptively adjust parameters for various traffic patterns based on contextual features, while also utilizing GAN to learn these traffic patterns from sparsely sampled data. Zhang et al. (Zhang et al., 2021c) models the intricate correlations between road traffic and urban conditions in their proposed conditional GAN-based traffic estimation model. Unlike them, the authors in (Yuan et al., 2023) introduced parameterization of noise factors in a diffusion-based model, transforming the traffic matrix estimation problem into a gradient-descent optimization task.
3.3. Traffic Data Mining
In this section, we introduce the related works that utilize generative AI techniques to analyze traffic data from various sources, including trajectories, traffic images, and traffic videos, as summarized in Table 4.
Trajectories. Trajectories refer to a sequence of timestamped positions that describe the movement or path of an object over time, which provides valuable insights for analyzing various aspects such as urban transportation modes and human decision-making preferences. To identify the transportation modes, the authors in (Zhang et al., 2022a) introduced a semi-supervised method based on the Dirichlet VAE, which fuses geographic information with motion features from GPS trajectories. Zhang et al. (Zhang et al., 2020d) proposed a trajectory GAIL framework to learn human decision-making behaviors. This framework modeled the human decision processes using Markov decision processes with varying lengths to capture long-term decision dependencies. Further, it inferred the decision-making strategy from the historical dataset of the human agent in an inverse manner.
Traffic images. In autonomous driving scenarios, traffic images captured by cameras offer vital information about the surrounding driving environment, enabling autonomous vehicles to make accurate decisions. To overcome the challenge of limited traffic image diversity in real urban environments, Li et al. (Li et al., 2020b) introduced a GAN-based framework with structural information. This approach enables various urban image transformations while maintaining the integrity of foreground objects and image structural details. Zhao et al. (Zhao et al., 2022a) focused on the quality of an in-vehicle image. They proposed a GAN-based single-image raindrop removal network that combines task-specific visual attention and content perception mechanisms, allowing the network to focus on raindrop regions and their surroundings for effective raindrop removal.
Unlike them, some researchers focus on extracting valuable information from traffic images for tasks such as vehicle detection (Wang et al., 2021b), license plate identification (Lee et al., 2018; Han et al., 2020a; Ibrahim et al., 2022; Boby et al., 2023), traffic density recognition (Gawali and Deore, 2023), driving behaviors (Wijaya et al., 2022) and map enrichment (Zhong et al., 2022). These tasks often face challenges related to low resolution and inadequate lighting conditions. For example, the authors in (Hassan Onim et al., 2023) presented a novel vehicle detection approach that utilizes two different GANs to address low-light traffic scenarios. Boby et al. (Boby et al., 2023) conducted a comparative analysis between a GAN-based super-resolution model and an iterative refinement approach utilizing diffusion-based super-resolution models for license plate detection. In addition, the authors in (Gawali and Deore, 2023) presented a dual-discriminator conditional GAN for recognizing traffic density in both homogeneous and heterogeneous traffic scenarios.
Traffic videos. Compared to traffic images, traffic videos provide a wealth of redundant traffic information for autonomous vehicles, guaranteeing both safety and efficiency in navigation. A pivotal stage in numerous intelligent video processing applications, such as vehicle behavior analysis, is the segmentation of moving objects. RMS-GAN (Patil et al., 2020) is an end-to-end GAN using a recurrent technique, which integrates foreground probability information using residual and weight-sharing techniques to achieve precise segmentation. Krishna et al. (Krishna and Denzler, 2014) progressively combined generative and discriminative models to efficiently classify activities in traffic videos. Cheng et al. (Cheng et al., 2021a) proposed a deep conditional generative model to detect interactions between vehicles and vulnerable road users. This model utilized a conditional VAE-based framework with Gaussian latent variables to capture the behavior of road users and predict interactions in a probabilistic manner.
Category Task Method Papers
Trajectory Transportation mode identification VAE (Zhang et al., 2022a)
Decision-making behavior analysis GAN (Zhang et al., 2020d)
Traffic image Image transformation GAN (Li et al., 2020b)
Image denoise GAN (Zhao et al., 2022a)
Image resolution enhancement GAN (Ibrahim et al., 2022; Boby et al., 2023)
Vehicle classification GAN (Wang et al., 2021b)
Vehicle detection GAN (Hassan Onim et al., 2023)
License plate recognition GAN (Lee et al., 2018; Han et al., 2020a)
Traffic density identification GAN (Gawali and Deore, 2023)
Behavior detection GAN (Wijaya et al., 2022)
Road annotation GPT (Juhász et al., 2023)
Flow inference Normalizing Flow (Zhong et al., 2022)
Traffic video Vehicle-road user interaction detection VAE (Cheng et al., 2021a)
Object segmentation GAN (Patil et al., 2020)
Activity classification GAN (Krishna and Denzler, 2014)
Table 4. Related works about traffic data mining.
3.4. Traffic Anomaly Detection
Traffic anomalies refer to unusual events, patterns, or behaviors within a transportation system. These anomalies can disrupt the regular traffic flow or pose potential safety risks. Numerous studies have explored the use of generative AI techniques for traffic anomaly detection (Qiu et al., 2019; Lin et al., 2020; Kumar et al., 2020; Xie et al., 2023). Chen et al. (Chen et al., 2021b) presented a multi-modal GAN model to detect and classify the traffic events, leveraging data from multiple modalities to enhance detection accuracy. The authors in (Santhosh et al., 2021) proposed a hybrid architecture that combines CNN and VAE to detect and classify anomalies. Different from the above GAN-based methods, the work by Kang et al. (Kang et al., 2022) employed conditional normalizing flow to detect abnormal temporal snapshots, followed by road segment-level anomaly detection using a kernel density estimator. A-VAE (Aslam and Kolekar, 2023) was designed as an attention-based VAE architecture for anomaly detection in traffic videos. It utilized a combination of 2D CNN and BiLSTM layers integrated with an attention network for effective representation learning.
Cyberattacks would pose significant safety risks in the context of automated driving. To address it, Zhao et al. (Zhao et al., 2022c) presented a GAN-based intrusion system for connected vehicles to evaluate the legitimacy of incoming messages. Ding et al. (Ding et al., 2019) investigated the susceptibility of deep generative models in autonomous driving to Trojan attacks, revealing that manipulating training data and injecting carefully crafted data can instill Trojan behaviors in a deep generative model without altering its original training objective. Li et al. (Li et al., 2022c) identified two potential cyberattack scenarios: individual vehicle-targeted attacks and sensor measurement data injection attacks. To address these threats, they introduced a real-time anomaly detection model using GANs for detection purposes.
4. Generative AI for Traffic Prediction
Traffic prediction refers to estimate or forecast future traffic conditions, encompassing various aspects such as travel demand, travel time, traffic flow, and the movements and behaviors of both vehicles and individuals within the transportation system. Traffic prediction is vital for urban traffic planning and management. However, achieving accurate predictions presents significant challenges. First, urban traffic systems exhibit complex spatial and temporal dynamics. Second, traffic conditions are highly dynamic, susceptible to rapid changes due to factors like real-time incidents, ongoing construction activities, and even human behavior. Third, obtaining high-quality data from various sensors, cameras, and GPS devices is important for traffic prediction, but such data is often limited and incomplete. Generative AI technology offers powerful solutions to overcome these challenges. In our comprehensive literature review, we concentrate on the utilization of generative AI in addressing traffic prediction challenges. We categorize these works into four main aspects: human-related, vehicle-related, road segment-related and region-related traffic prediction, encompassing the key components of transportation systems. Their relation is given in Figure 4.
Figure 4. The relation between different topics in generative AI for traffic prediction research.
4.1. Human-related Traffic Prediction
Human-related traffic prediction involves forecasting individual behavior and human mobility.
Individual behavior. We discuss individual behavior prediction based on their roles within transportation systems. For drivers, modeling and predicting driver behavior plays a crucial role in enhancing safety and preventing risky actions (Ivanovic et al., 2020; Bao et al., 2021). For example, Bao et al. (Bao et al., 2021) introduced a probabilistic sequence-to-sequence approach utilizing a conditional VAE to predict various driving behaviors. For pedestrians, accurately forecasting human motion behavior is vital for autonomous vehicles to plan preemptive actions and avoid collisions, ensuring pedestrian safety (Sun et al., 2019; Cui et al., 2021). For instance, in (Cui et al., 2021), the authors proposed a temporal convolutional GAN for high-fidelity future pose prediction, incorporating two discriminators: one for fidelity assessment and another for consistency evaluation in long sequences.
Human mobility. Human mobility refers to an individual’s movement from one location to another. Feng et al. (Feng et al., 2023b) employed GAIL to model courier decision-making in route choice, enabling the prediction of their future routes. The COVID-19 pandemic presented a great challenge in predicting human mobility. Bao et al. (Bao et al., 2020) introduced a spatio-temporal conditional GAN Covid-GAN to estimate mobility during the pandemic. Further, Covid-GAN+ (Bao et al., 2022) addressed spatial heterogeneity by using the local Moran statistic and outliers by redesigning the training objective to learn the estimated mobility changes.
4.2. Vehicle-related Traffic Prediction
In the reviewed literature, vehicle-related traffic prediction tasks encompass vehicle trajectory prediction and crash risk prediction.
Vehicle trajectory. A vehicle trajectory refers to the path a vehicle takes from one location to another over a specific time period. Precise vehicle trajectory prediction is vital for a variety of traffic-related applications, including traffic management, navigation, and route planning. However, achieving precise trajectory predictions is a challenging task for several reasons. First, vehicle movements are influenced by various factors, including driver behavior, traffic conditions, road infrastructure, and unexpected events. Second, traffic conditions can change rapidly due to incidents, congestion, or accidents. Third, the accuracy of trajectory prediction models heavily relies on the quality and quantity of input data, such as GPS data. Noisy or incomplete data can lead to inaccurate predictions. To address these challenges, more researchers proposed different methods based on generative AI technology to predict vehicle trajectory.
Most problems of vehicle trajectory prediction are defined as using historical trajectories to infer future trajectories. Formally, given a vehicle $i$ with a historical trajectory $\mathbf{x}_{i}=\{r_{i}^{t}\},t=1,2,...,T$, where $r_{i}^{t}$ is location information (e.g., longitude and latitude) at time $t$. The task is to learn a mapping function $g(\cdot)$ to predict the future trajectory $\mathbf{y}_{i}=\{r_{i}^{t}\},t=T+1,...,T+n$. Many scholars adopt different generative AI techniques to address above challenges.
VAE. Several researchers have designed prediction models based on VAE (Cheng et al., 2020; Kim et al., 2021; Xing et al., 2022; Oh and Peng, 2022). DVAE (Neumeier et al., 2021) incorporated expert knowledge into the decoder of a VAE, which can provide the similar prediction accuracy and enhance interpretability. The authors in (Jagadish et al., 2021, 2022) utilized conditional VAEs to encode past trajectories of vehicles and traffic scenes for vehicle path prediction. Gui et al. (Gui et al., 2022) incorporated more information like yaw angle and velocity in their proposed conditional VAE-based trajectory prediction model.
GAN. GAN is also an alternative technique for trajectory prediction (Roy et al., 2019; Li et al., 2021a; Zhou et al., 2021b). For example, Kang et al. (Kang et al., 2020a) applied social GAN, originally designed for pedestrian trajectory prediction, for the prediction of vehicle trajectories. Differently, Zhou et al. (Zhou et al., 2021b) focused on predicting trajectories of surrounding vehicles. They introduced a self-attention social GAN, which not only addressed the challenge of missing information from long trajectory sequences but also better modeled the interactions between vehicles.
GPT. Recently, GPT models has great potential in vehicle trajectory prediction (Su et al., 2022; Feng et al., 2023a; Vishnu et al., 2023). Su et al. (Su et al., 2022) proposed a crossmodal transformer based generative framework, which leveraged sequences of cues from various modalities and pedestrian attributes for more accurate predictions. In this model, transformer is adopted to capture the cross-relation patterns of different modality pairs. Feng et al. (Feng et al., 2023a) developed a pre-trained large traffic model using the Transformer architecture. This model has ability to capture the diverse trajectories within a population of vehicles, leading to higher accuracy in vehicle trajectory prediction.
DPM. Diffusion models also attract more researchers’ attentions. For instance, EquiDiff (Chen et al., 2023a) combined the conditional diffusion model with an equivariant transformer, leveraging the geometric properties of position coordinates and social interactions among vehicles for accurate prediction.
Crash risk. Crash risk prediction is important in traffic safety management and precaution. Due to imbalanced data between crash and non-crash cases, accurately predicting crash risk is challenging. To deal with this, Cai et al. (Cai et al., 2020) proposed a deep convolutional GAN model to generate more synthetic data about crashes. Similarly, in (Man et al., 2022), Wasserstein GAN was developed to address the imbalanced data. Different from approaches using GAN, Zhang et al. (Zhang et al., 2022d) proposed an enhanced accelerated testing method based on Normalizing Flows. This method leverages Normalizing Flows to learn the distribution of rare events and preserve the correlated relationships between scenario variables.
4.3. Road Segment-related Traffic Prediction
Road segment-related traffic prediction focuses on forecasting traffic conditions (e.g., traffic flow, speed, travel time) for specific segments or sections of roads within a transportation network. Accurate prediction faces three main challenges. First, traffic conditions within the same road segment can vary over time, but there is often similarity between adjacent time periods. Additionally, the traffic state of one road segment can affect neighboring segments, resulting in complex spatial and temporal patterns in traffic conditions. Second, the limited number of road sensors and the possibility of sensor malfunctions can result in missing or inaccurate data, posing challenges for ensuring data quality and completeness. Third, external factors like accidents and weather further enhance the uncertainty of prediction. Generative AI technology has gained significant attention from scholars due to its ability to learn data distributions and model uncertainties. Numerous models based on generative AI technology are proposed for predicting traffic flow and travel time.
Traffic flow prediction. The task of traffic flow prediction is to learn a non-linear function $h(\cdot)$ from previous $T$ time steps of traffic conditions to forecast traffic conditions for the next $T^{{}^{\prime}}$ steps based on road network graph $G$. Many scholars employ different generative AI techniques to tackle this task. We summarize the related works in Table 5.
Task Method Spatio-temporal modeling Papers
Traffic flow prediction VAE Temporal: LSTM (Jin et al., 2022a)
GAN Temporal: LSTM (Xu et al., 2020a)
Spatial: GCN (Yu et al., 2020; Zheng et al., 2022)
Spatial: GCN, Temporal: RNN (Jin et al., 2022b)
Spatial: GCN, Temporal: GRU and self-attention (Khaled et al., 2022)
Spatial: GCN, Temporal: TCN (Li et al., 2022b)
Saptial: GraphSAGE; Temporal: LSTM (Zhao et al., 2022b)
Spatial: GCN; Temporal: attention network (Devadhas Sujakumari and Dassan, 2023)
Normalizing flows Autoregressive network with convolution operations (Zand et al., 2023)
DPM Spatial: GCN, Temporal: TCN (Wen et al., 2023)
Temporal: RNN (Rasul et al., 2021)
Travel time estimation GAN Spatial: GCN (Song et al., 2021)
DPM Masked Vision Transformer (Lin et al., 2023b)
Table 5. Related works about road segment-related traffic prediction.
VAE. Some researches propose VAE-based models. The authors in (Boquet et al., 2020) designed a VAE-based prediction model with the ability to learn the data distribution. PFVAE (Jin et al., 2022a) employed LSTM as the autoencoder and designed the VAE as a predictor for time-series data to deal with the noise issues.
GAN. GAN-based models are commonly adopted in this task (Lin et al., 2018; Zhang et al., 2019b; Saxena and Cao, 2019; Zhang et al., 2021e; Devadhas Sujakumari and Dassan, 2023). For example, ForGAN (Koochali et al., 2019) employed GAN to learn the data generating distribution and generate probabilistic traffic forecasts. Furthermore, many researchers designed several methods to model spatial and temporal correlations in their proposed GAN-based prediction model (Zang et al., 2019; Xu et al., 2020a; Yu et al., 2020; Zheng et al., 2022; Jin et al., 2022b; Khaled et al., 2022; Li et al., 2022b; Zhao et al., 2022b). For example, PL-WGAN (Jin et al., 2022b) utilized GCN, RNN and attention mechanism to model spatiotemporal correlations in a Wasserstein GAN-based model. ASTGAN (Liu and Zhang, 2022) used the attention mechanism and an mask graph convolutional recurrent network to characterize the temporal and spatial dependencies.
Normalizing flows. Consider the capability of conditional normalizing flows in representing complex distributions characterized by high dimensionality and strong interdimensional relationships. MotionFlow (Zand et al., 2023) utilized conditional normalizing flows for traffic prediction, which conditionally models the output distributions on spatio-temporal input features in an autoregressive manner.
DPM. Because DPM can effectively handle uncertainties within spatio-temporal graph neural network (STGNN), DiffSTG (Wen et al., 2023) combined the spatiotemporal learning capability of STGNN with the uncertainty measurement of DPM, resulting in higher prediction accuracy. The authors in (Rasul et al., 2021) developed an autoregressive denoising diffusion model for multivariate probabilistic time series forecasting, utilizing gradient estimation to draw samples from the data distribution at every time step.
Travel time estimation. Travel time estimation task aims to estimate the duration it takes to travel between two locations or along a specific route given the departure time. Several scholars employ generative AI techniques to address this task (Song et al., 2021; Lin et al., 2023b). To be specific, GCGTTE (Song et al., 2021) integrated GCN and GAN to estimate the travel time in a probabilistic distribution form. Lin et al. (Lin et al., 2023b) proposed a two-stage diffusion-based origin-destination (OD) travel time estimation model. They incorporated a conditioned pixelated trajectory denoiser into the model to capture correlations between OD pairs and historical trajectories.
4.4. Region-specific Traffic Prediction
Region-specific traffic prediction focuses on forecasting traffic conditions such as traffic flow within specific geographic regions in a transportation network. The main challenges in achieving accurate predictions include complex spatio-temporal dependencies, limited data, and noise. Recent works (Li et al., 2020a; Zhang et al., 2020c, 2019a; Naji et al., 2021; Zhang et al., 2023a; Feng et al., 2021) have aimed to employ generative AI techniques to tackle these challenges. To be specific, Zhang et al. (Zhang et al., 2022b, 2023a) proposed a spatiotemporal GAN model for predicting traffic flow in subway stations, which used gated TCN and weight sharing GCN to capture spatiotemporal correlations. Kang et al. (Kang et al., 2020b) incorporated gated convolution and dilated convolution to capture the traffic patterns in urban regions for region-level traffic prediction. Li et al. (Li et al., 2022a) proposed a seq2seq spatial-temporal semantic GAN model, which takes into account both semantic factors and spatio-temporal features. The authors in (Naji et al., 2021) incorporated recurrent network mode and conventional network model into GAN to predict the taxi demand in specific areas.
5. Generative AI for Traffic Simulation
Traffic simulation involves the mathematical modeling of transportation systems, which can be used to generate the movement and behavior of vehicles and pedestrians. This technology can be applied in various fields, including urban planning, evaluation of policy changes, and testing and validating autonomous vehicle algorithms. However, achieving a more realistic traffic simulation poses several challenges. First, real-world traffic scene data is typically collected through road cameras or sensors. However, obtaining such data is not only costly but also challenging to capture in rare or hazardous situations. Second, traffic dynamics are inherently complex due to the numerous variables involved, ranging from individual driver behaviors, decisions, and reactions to the unpredictable nature of external factors such as weather or sudden events. Capturing the intricacies of these dynamics is a considerable challenge. Additionally, the interconnected relationships between vehicles, infrastructure, and pedestrians further complicate the situation. Existing works using traditional deep learning methods (Halim et al., 2016; Xiao et al., 2022) often require extensive labeled data, and may not generalize well to unseen or rare situations. In addition, these methods may struggle to capture the vast array of dynamic interactions in traffic scenarios (Wei et al., 2014; Li et al., 2021b).
Generative AI techniques offer potential solutions for traffic simulation challenges. They can create realistic traffic scenarios, including rare or extreme situations that might be missing from real-world data. Their adaptability allows them to adjust to ever-changing traffic dynamics, thereby capturing intricate interactions more holistically. Furthermore, by learning from a small number of real traffic data, they can refine their simulations continuously, ensuring they mirror the complexities of real-world traffic more closely. Next, we will introduce the related works using generative AI techniques from three aspects: driver behavior simulation, traffic scenario generation and traffic flow generation. We draw their relation in Figure 5.
Figure 5. The relation between different topics in generative AI for traffic simulation research.
5.1. Driver Behavior Simulation
In autonomous driving scenarios, the focus is on how drivers make decisions in response to traffic conditions (Ghosh et al., 2016; Yun et al., 2019; Chen et al., 2022b; Ye et al., 2022; Singh et al., 2023). Yun et al. (Yun et al., 2019) utilized GAN to learn the human-driving abilities in complex traffic scenarios. Chen et al. (Chen et al., 2022b) proposed a novel two-stage driver model that integrates model-based and data-driven techniques, leveraging reward-augmented GAIL to produce reliable human driving behaviors. Jin et al. (Jin et al., 2023) proposed a generative driver agent simulation framework built upon LLM to simulate human driving behaviors based on human driver experience. To alleviate the computational load of large-scale simulations, Ye et al. (Ye et al., 2022) proposed an efficient method for dynamically calculating agent decision parameters in ABM calibration by using VAEs to accelerate state transfer probability computation.
Lane changing and car following are two common driving behaviors, and numerous studies have explored methods for simulating these behaviors (Lin et al., 2021; Ma et al., 2023; Li et al., 2023b; Dong et al., 2023). Ma et al. (Ma et al., 2023) introduced a physics-informed conditional GAN that integrates the advantages of both physics-based and deep learning models to to improve the modeling of car-following behavior over multiple time steps in complex traffic flow environments. Lin et al  (Lin et al., 2021) introduced a constrained-GAIL framework that employs reward augmentation to guide driving agents in avoiding undesirable states by imposing additional constraints through a manually designed reward function. In (Dong et al., 2023), the authors leveraged the transformer and diffusion models to generate lane-changing trajectories that closely mimic human behavior.
Unlike autonomous driving scenarios, in transportation services, a variety of factors such as estimated travel demand tend to dominate drivers’ behaviors. For example, Zhang et al. (Zhang et al., 2020a) developed conditional GAIL model capable of learning the driver’s decision-making tendencies and strategies through knowledge transfer among taxi driver agents and across various locations.
5.2. Traffic Scenario Generation
Task Method Controllability Papers
Vehicle trajectory generation VAE Unconditional (Chen et al., 2021c; Zhang et al., 2022e)
GAN Unconditional (Kumar et al., 2023; Liao et al., 2022; Ozturk et al., 2020; Liao et al., 2022; Choi et al., 2021)
DPM Unconditional (Sun et al., 2023; Zhu et al., 2023; Xu et al., 2023d, b)
VAE, GAN Unconditional (Krajewski et al., 2018, 2019)
GAN Conditional (Arnelid et al., 2019)
DPM Conditional (Zhong et al., 2023a, b)
Human trajectory generation GAN Unconditional (Xiong et al., 2023)
DPM Conditional (Rempe et al., 2023)
Anomaly data generation VAE Unconditional (Cai et al., 2020; Islam et al., 2021)
GAN Unconditional (Chen et al., 2021e, 2022a; Huo et al., 2021)
GAN Conditional (Zarei and Hellinga, 2021)
Table 6. Related works about traffic scenario generation.
Traffic scenario generation refers to produce specific traffic situations or anomalous events for the purpose of testing and validating autonomous driving systems. As autonomous vehicles need to operate safely in a vast array of driving conditions, it is crucial for developers to test these systems against a wide range of potential real-world scenarios, especially those that are rare or potentially hazardous. In this literature review, we will focus on three key areas: generating vehicle trajectories, generating human trajectories, and generating anomaly data, as summarized in Table 6.
Vehicle trajectory generation. Many researchers use generative AI technology to generate more vehicle trajectory data (Krajewski et al., 2018, 2019; Choi et al., 2021; Liao et al., 2022; Zhang et al., 2022e; Chen et al., 2021c; Ozturk et al., 2020; Kumar et al., 2023; Sun et al., 2023; Xu et al., 2023d). In (Krajewski et al., 2018), the authors introduced trajectory GAN and trajectory VAE to generate realistic synthetic lane change maneuver trajectories without expert knowledge. Diff-Traj (Zhu et al., 2023) was proposed for generating GPS trajectories by effectively learning spatio-temporal features from historical trajectories and leveraging the generation capability of the diffusion model. Xu et al. (Xu et al., 2023b) designed a generative AI-based autonomous driving architecture that utilizes large text-to-image models based on DPM to generate conditioned traffic and driving data in simulations. Differently, some searchers focused on the controlled vehicle trajectory generation (Arnelid et al., 2019; Zhong et al., 2023a, b). For example, the authors in (Arnelid et al., 2019) introduced a recurrent conditional GAN model that generates sensor errors with long-term temporal correlation in both the generator and the discriminator networks. Zhong et al. (Zhong et al., 2023a) introduced a scene-level conditional diffusion model that incorporates a spatio-temporal transformer architecture to produce both realistic and controllable traffic data.
Human trajectory generation. Regarding human trajectory generation, Xiong et al. (Xiong et al., 2023) proposed a semantic-guiding GAN model equipped with semantic knowledge to generate human trajectories. Rempe et al. (Rempe et al., 2023) studied controlled pedestrian trajectory generation. They utilized guided diffusion modeling for test-time controllability, and integrating it with a physics-based humanoid controller for a closed-loop system that places large crowds in various terrains while meeting user-defined goals.
Anomaly data generation. Anomaly data such as traffic accident and crash data is essential in autonomous driving for safety assessment and scenario testing. However, this type of traffic data is hard to collect. For example, as illustrated in (Cai et al., 2020), the ratio of crash data to non-crash data is 1 to 11,000, indicating a significant imbalance. Several works use generative AI techniques to address this problem (Islam et al., 2021; Chen et al., 2021e, 2022a; Huo et al., 2021; Zarei and Hellinga, 2021). For instance, Islam et al. (Islam et al., 2021), employed VAE to encode events into a latent space, from which they subsequently sampled to generate crash data. In (Huo et al., 2021), a text-to-traffic GAN model was proposed, which integrates traffic data with semantic information collected from social media to produce traffic situations. Zarei et al. (Zarei and Hellinga, 2021) developed a conditional GAN-based model that utilizes crash count as a condition to generate crash data.
5.3. Traffic Flow Generation
Traffic flow generation refers to generate vehicular traffic volumes or crowd flow within a specific area or between regions in a transportation network. Numerous works adopt generative AI techniques to address this task (Chen et al., 2021a; Li et al., 2022d; Zhou et al., 2023; Rong et al., 2023a, b). Specifically, Li et al. (Li et al., 2022d) presented an attentive dual-head spatial-temporal GAN model for generating crowd flow. This model incorporated an attentive mechanism for tracking temporal changes and a self-attention network to capture intricate spatiotemporal crowd flow patterns, and it adopted a dual-head discriminator that implements a training strategy with two distinct objectives to mitigate the adverse effects of rapid overfitting. Zhou et al. (Zhou et al., 2023) constructed an urban knowledge graph to extract region features and model the urban environment, and they subsequently designed a knowledge-aware spatio-temporal diffusion model for generating urban flow. In (Rong et al., 2023a), the authors employed a graph denoising diffusion technique to generate urban OD flow, enabling the learning of how the likelihood of nodes and edges are interconnected in the OD network.
6. Generative AI for Traffic Decision-making
In autonomous driving scenarios, traffic decision-making is a series of decision-making processes made by vehicles based on current dynamic traffic environment and task requirements when driving on the road, such as driving routes, speeds, driving modes and so on. It is a challenging task for ensuring that vehicles make safe and accurate decisions in complex environments. The first challenge we face is dealing with the dynamic and unpredictable nature of traffic environments. These environments are constantly changing, with various factors like road conditions, unexpected moves by human drivers, and sudden decisions by pedestrians and cyclists. For example, a driver might suddenly brake or change lanes without warning, or pedestrians might cross the road unexpectedly. These uncertainties compel autonomous systems to be perpetually alert and adaptive, which becomes computationally intensive and requires robust prediction algorithms. The second challenge involves rare but critical events that can test the safety of autonomous vehicles, such as a child running onto the road or a mechanical failure causing loss of control. Gathering real-world data on these rare events is difficult but crucial for creating a comprehensive training dataset. Without exposure to these scenarios, autonomous systems may not know how to handle them properly, potentially leading to safety risks. The third challenge relates to the limitations of sensors like LiDAR, RADAR, and cameras. These sensors have restrictions in terms of their range, resolution, and susceptibility to environmental factors like fog, rain, and snow. These limitations can sometimes lead to uncertainty or incorrect decision-making by the autonomous vehicle, which can compromise both safety and efficiency.
Traditional decision-making methods, such as rule-based methods (McNaughton et al., 2011; De Beaucorps et al., 2017) and game theory-based methods (Wang et al., 2015; Hang et al., 2020; Zhao et al., 2021a), often have limited effectiveness in dealing with complex traffic situations, while reinforcement learning (RL), as a flexible, efficient, and powerful method, has gained ground in autonomous driving decision-making. However, its general application is hindered by training complexity, especially in handling multiple tasks. Generative AI technology holds promise in tackling these challenges. Leveraging their robust generative capabilities, they can generate realistic training scenarios mimicking the unpredictability of real-world driving, reducing the need for extensive road testing. Moreover, they can produce rare traffic events to ensure autonomous systems are robust against critical scenarios. Additionally, generative AI can enhance sensor data, helping vehicles adapt to challenging environmental conditions.
Recently, researchers have shifted their focus to using generative AI techniques in traffic decison-making. Wang et al. (Wang et al., 2022) proposed an autonomous decision-making approach for high-speed trains, integrating GAN-based learning approach and a distributed tracking controller. Addressing data insufficiency, they introduced a GAN-based data augmentation method to generate time series samples for training purposes. They then developed a trajectory prediction network that compute the reference speed trajectory in real-time conditions. Further, they designed a distributed tracking control model, utilizing model predictive control and dual decomposition, to optimize train movement.
Liu et al. (Liu et al., 2023a) proposed a multi-task decision-making GPT model for autonomous driving at unsignalized intersections. By leveraging the exceptional learning capabilities and computational efficiency of RL, this model had ability to simultaneously manages multiple driving tasks. The authors first trained RL expert models for single decision-making tasks, and then used the sampled expert data to guide the offline GPT training. As a result, the model outperforms conventional single-task RL decision-making models.
7. Open Challenges and Future Directions
While generative AI methods have demonstrated much success in the realm of intelligent transportation systems, several challenges must be addressed. In this section, we will discuss the primary open challenges, followed by an introduction to some potential research directions.
7.1. Open Challenges
In this subsection, we will discuss several open challenges for generative AI methods in intelligent transportation systems.
Multi-modal traffic data. Multi-modal traffic data integrates various data sources and types, including roadside camera videos, traffic signal sensor data, weather conditions, traffic incident reports, and vehicle GPS trajectories. These data provides a comprehensive view of dynamic traffic conditions in transportation systems. Leveraging generative AI technology to address traffic-related problems with multi-modal data present several challenges. Firstly, the correlations of multi-modal data are more complex than those of single-modal data. For example, the characteristics of images and texts, as well as the differences in data distributions and representation methods, are distinct. Secondly, accurate alignment between different modalities is required, e.g., matching traffic camera images with traffic sensor data or weather information at the same time. Additionally, training models on multi-modal data can be more unstable, especially when maximizing mutual information across modalities. Therefore, how to effectively use generative AI technology to tackle transportation challenges using multi-modal data remains challenging in the field of transportation.
Complex spatio-temporal correlations. Traffic involves complex relationships in time, space, and spatiotemporal interactions. In the temporal dimension, traffic data can be influenced by seasonal variations such as weekdays, weekends, and holidays. Simultaneously, the traffic flow at a specific moment can be impacted by the traffic patterns at that location over the preceding few hours. In the spatial dimension, traffic data involves the traffic conditions at different locations, and it is necessary to consider the traffic flow and congestion between different locations, and analyze their correlation. Moreover, there is a complex interaction between time and space in traffic conditions, and the traffic flow at a location may be affected by the traffic flow at adjacent locations in the past few hours. While prior research has explored the use of RNNs, CNNs, and GNNs in generative AI-related solutions to model spatiotemporal relationships, yielding some promising results, most of these approaches focus on local or micro-scale spatiotemporal modeling and fail to capture global or macro-scale spatiotemporal patterns. Furthermore, spatiotemporal modeling remains somewhat disconnected from the core principles of generative AI. Thus, the challenge in leveraging generative AI technology to address traffic-related issues lies in effectively modeling the intricate spatiotemporal correlations within traffic data.
Sparse or missing traffic data. The sparsity or missing of traffic data can be attributed to various factors, including equipment failures or routine maintenance, network communication issues, and concerns related to the costs and privacy of data collection. The sparse or missing nature of traffic data can have significant implications for traffic management and planning, including inaccurate traffic predictions and an incomplete understanding of traffic conditions. Generative AI has powerful data generation capabilities and can generate relatively realistic data by leveraging the characteristics of existing data. To address the problem of data sparsity or absence, many researchers use generative AI technology to perform tasks such as data imputation, traffic data estimation, and traffic scene generation, but there are still shortcomings. First, due to the spatial-temporal complexity of traffic data, generative AI models may generate some seemingly realistic data that is inconsistent in space and time. Second, when generating data, the generative model may tend to generate common patterns in training data while neglecting less common cases. This may lead to a lack of diversity in the generated data and inability to cover all the changes in traffic data. Third, different degrees of sparsity also have a significant impact on the generation quality, and how to flexibly cope with this is an important issue. Therefore, overcoming the challenges of data sparsity and absence in addressing traffic issues with generative AI technology remains an important problem that needs further research.
Adversarial attacks. In the realm of intelligent transportation, generative AI technology offers cutting-edge solutions that drive continuous advancements in traffic management and optimization. However, the widespread adoption of this technology also introduces new security risks, particularly those associated with adversarial attacks (Foo et al., 2023). One of the most notable concerns is the backdoor attack, where malicious actors may inject harmful data during the model training phase, causing the model to generate inappropriate outputs under specific conditions. In complex intelligent transportation networks, a single misjudgment or erroneous decision can trigger a chain reaction of traffic accidents or congestion, resulting in substantial economic and societal losses. For instance, an intelligent transportation monitoring system compromised by a backdoor attack may erroneously classify a safe vehicle as a potential threat, or an autonomous vehicle may make incorrect driving decisions as a result. Therefore, the critical challenge lies in accurately identifying potential risk points and devising effective methods to ensure the efficient and secure operation of transportation systems.
Model interpretability. Model interpretability is crucial in intelligent transportation systems. For example, in autonomous driving scenarios, it is necessary to predict the behavior of other road users (pedestrians, cyclists, other vehicles) to ensure safe driving. This requires the model to explain why a particular behavior is considered the most likely, such as why the model believes a pedestrian in front may cross the road. Although generative AI models perform well in processing traffic data and optimizing transportation systems, these models are often considered black boxes, making it difficult to understand their decision-making processes and reasoning foundations. However, designing interpretable generative AI models is challenging. This is because traffic systems are very complex, involving many variables and uncertain factors such as road conditions, traffic flow, pedestrian behavior, and other vehicle behavior. These complexities and uncertainties make it difficult to design a simple and accurate model to explain traffic data and autonomous driving decisions. In autonomous driving, autonomous driving decisions involve multiple levels, including perception, planning, and control, and designing an interpretable model requires being able to explain the decision-making processes at these different levels, which are often very complex.
Real-time requirements. In intelligent transportation systems, taking timely and effective response measures is of paramount importance. This becomes particularly critical when dealing with rapidly changing situations and unforeseen events. In such scenarios, traffic-related models must possess real-time perception capabilities, including the detection of other vehicles, pedestrians, and obstacles, while simultaneously making split-second decisions to ensure safe driving. In addition, traffic congestion can lead to rapid changes in traffic flow, so these models need to have the ability to predict real-time changes in traffic flow to take measures in advance and avoid traffic accidents and congestion. However, achieving this high level of responsiveness poses some challenges in generative AI-based solutions. First, real-time data may be affected by noise and incompleteness, so models must have the ability to handle imperfect data to accurately perceive the traffic environment. Second, certain generative AI methods may be very complex and require more time for reasoning. In real-time applications, it is necessary to balance the complexity and real-time performance of the model. Third, real-time applications should be equipped to handle exceptional circumstances like sensor failures or sudden events. Generative AI methods need to be capable of identifying such anomalies and executing appropriate actions to ensure both safety and reliability.
7.2. Potential Research Directions
Large-scale language models designed using generative AI technology have achieved significant success in fields such as NLP and image processing, which also provide crucial opportunities and research directions for addressing traffic-related challenges. In this subsection, we will introduce several important research directions from three distinct perspectives concerning large-scale traffic models.
Integrating multi-modal data for large-scale traffic models. Traffic data usually come from multiple different sources, including images and video data from traffic cameras, as well as GPS trajectory data from vehicles, covering a variety of different data types. These data offer multiple perspectives on the spatiotemporal dynamics of traffic conditions, facilitating a deeper comprehension of traffic operation’s underlying mechanisms and patterns. Therefore, these data serve as an important foundation for training large-scale models in the field of transportation. Efficiently using these data for model training is crucial and also an important research direction. In addition to aligning the data in the spatiotemporal dimension, it is necessary to unify their representation, mapping different data modalities to the same data space. This requires designing appropriate mapping methods to ensure that the mapped data can effectively capture spatial and temporal dependencies across different data modalities.
Designing and training well-performed large-scale traffic models. Traditional transportation models and algorithms are typically developed separately for specific tasks, necessitating significant time and resource investments for model adjustments when encountering new tasks or changing problem requirements. Furthermore, these conventional models often exhibit excessive specialization, rendering cross-task transfer learning challenging and thereby limiting their overall generalizability and scalability. In contrast, large-scale language models based on generative AI technology have higher generalizability and scalability, and possess the potential for transfer learning in a variety of traffic tasks. For instance, in autonomous driving, which encompasses continuous tasks such as perception, prediction, and decision-making, we can design a large-scale transportation model to handle all three tasks simultaneously. In model design, special attention needs to be paid to shared features, such as spatiotemporal characteristics, and these features should be modeled effectively while combining the uniqueness of each task to achieve consistent excellent performance across all tasks. Additionally, another potential research direction is to develop large-scale traffic models driven by both data and knowledge, leveraging the diversity of traffic data and the reliability of traffic physical knowledge to effectively model traffic patterns. Finally, it is crucial to consider how to ensure that the designed large-scale model can achieve the desired performance and effectiveness. This requires the development of reasonable dataset annotation strategies and training strategies.
Traffic planning and decision-making based on large-scale language models. Traffic planning and decision-making are important research problems in the transportation domain. In real-world scenarios, traffic conditions are constantly changing, with the possibility of incidents such as traffic congestion, accidents, and adverse weather conditions, all of which may impact the effectiveness of traffic planning and decision-making. Therefore, it is essential to develop real-time traffic planning and decision-making algorithms based on large-scale language models. This entails exploring how to combine the real-time traffic perception capabilities of generative AI technology with the decision-making capabilities of RL. Such an approach holds the potential to yield more efficient and adaptable solutions to tackle the intricacies and uncertainties inherent in traffic management. As an illustrative example, consider driving decision-making in autonomous vehicles. Vehicles need to make decisions such as overtaking, decelerating, and stopping based on perceived data and environmental conditions. Large-scale models, equipped with powerful knowledge reasoning and generation capabilities, are expected to achieve precise traffic decisions in complex traffic environments. Therefore, future research can focus on how to design algorithms based on large-scale language models to make accurate decisions while ensuring a balance in real-time performance and safety.
8. Conclusions
Generative AI technology is playing an increasingly important role in perception, prediction, simulation, and control of transportation systems, providing powerful tools and methods to enhance traffic efficiency, safety, and sustainability. In this survey, we systematically investigate the critical solutions of generative AI techniques within intelligent transportation systems, and explore the challenging issues that generative AI has not yet fully addressed in this field. We hope that this survey will promote further innovation and development of generative AI technology in the field of intelligent transportation systems, assist researchers in effectively utilizing generative AI technology to tackle more practical traffic problems, and encourage the development of more advanced generative AI methods targeting the transportation domain.